# DeepAndes: A Self-Supervised Vision Foundation Model for Multi-Spectral Remote Sensing Imagery of the Andes

### v1

@@ IV Data and Experiments
IV-A Data Pre-processing
Firstly, the raw satellite imagery is pre-processed into digitized archaeological images for the computer system in preparation for training deep learning models. The remote sensing satellite images used in this work are collected by the WV-2 and WV-3 satellites provided by the Digital Globe Foundation, following color correction and orthographic correction using a coarse digital elevation model (DEM). The data are then pan-sharpened using the Bayesian fusion algorithm from Orfeo-Toolbox [21] to increase the spatial resolution of the multi-spectral imagery to 0.31 m for WV-2 imagery to match the native resolution of WV-3 imagery. In this work, all 8 spectral bands (four standard colors—red, green, blue, and near-infrared 1—and four new bands—coastal, yellow, red edge, and near-infrared 2) are used. Lastly, the imagery is re-sampled from 32 bits to 8 bits to reduce storage size and computational load.
@@ IV-B SSL with Different Pre-training Scales
As introduced in Section III-A, after data pre-processing, we construct the self-supervised pre-training dataset sourced from the entire Peruvian Andes. In total, the imagery covers survey regions of approximately 488,640 $km^{2}$ and includes 8 distinct land cover types. We densely sample 3 million image patches of size 256 $\times$ 256 from all surveyed regions to ensure a diverse, large-scale pre-training dataset. In this work, we perform SSL on pre-training datasets of varying scales to robustly evaluate the impact of pre-training on downstream tasks. Specifically, the ViT-L backbone is pre-trained using 30K, 300K, and 3 million image patches, and the corresponding pre-trained models are then evaluated respectively. The same procedure is followed to create these pre-training datasets, with the only difference being the number of patches sampled from each land cover type.
@@ IV-C1 Imbalanced Loci Classification
Evaluation Metric. The F1 score is used as the evaluation metric for this imbalanced classification task, as it combines both sensitivity and precision,
$F1\ score\ =2\ \ast\ \frac{Precision\ \ast\ Sensitivity}{Precision\ +\ Sensitivity}$
where the Sensitivity (or Recall) determines the accuracy of the minority class classification and Precision indicates the probability of its correct detection. Particularly, the Precision is calculated as the ratio of true positives (TP) to the total number of positive predictions, which is the sum of true positives and false positives (FP):
@@
The Recall is the ratio of TP to the total number of actual positive samples, which is the sum of true positives and false negatives (FN):
@@ IV-C2 Image Instance Retrieval
$Precision@i\ =\frac{\text{Number of images retrieved at rank }i}{i}$
@@
IV-C3 Few-shot Loci Segmentation
To evaluate the foundation model’s few-shot learning performance on pixel-level feature recognition, we focus on the semantic segmentation of three specific types of loci (i.e., active buildings, active corrals, and archaeological corrals) in this study. Three small-scale binary semantic segmentation datasets are constructed: the Active Buildings dataset contains 48 images, the Active Corrals dataset contains 55 images, and the Archaeological Corrals dataset contains 46 images. K-fold cross-validation (K=5) is employed for robust evaluation. The segmentation head is simple and directly utilizes the learned features from pre-training. Specifically, patch-level features are extracted from the images using patch tokens and concatenated with a simple one-layer linear segmentation head. Prior to the five-fold cross-validation experiments, we conducted several random seed trials to identify optimized hyperparameters for model training. Similar to the few-shot loci classification, we also evaluate the pre-trained model’s segmentation performance at different downstream dataset scales, including 10, 20, and 30 images.
@@ IV-C3 Few-shot Loci Segmentation
  • $TP$ represents the number of pixels correctly classified as the loci in both the predicted and ground truth masks.
  • $FP$ represents the number of pixels incorrectly classified as loci in the predicted mask but not in the ground truth mask.
  • $FN$ represents the number of pixels incorrectly classified as background in the predicted mask but actually belonging to the loci in the ground truth mask.
@@ V-A Imbalanced Loci Classification
Both Table I and Figure 5 display the five-fold cross-validation results for imbalanced loci classification. Specifically, Table I details the mean F1 scores, Precision (P), and Recall (R). Figure 5a illustrates the Precision-Recall (PR) curves, while Figure 5b presents the confusion matrices for each model, highlighted using a representative hold-out fold.
As shown in Table I, the Scratch model struggles with limited and imbalanced data, converging only at $N_{\text{train}}$ = 729 (F1 = 0.544) and $N_{\text{train}}$ = 365 (F1 = 0.402), while failing entirely at smaller datasets. In contrast, DeepAndes models with SSL pre-training improve consistently across all dataset sizes. At $N_{\text{train}}$ = 72, FM3M achieves F1 = 0.83, Recall = 0.825, and Precision = 0.837, effectively balancing false positives and false negatives. Smaller pre-trained models decline in performance, with FM300K dropping to F1 = 0.728, R = 0.671, P = 0.812 and FM30K to F1 = 0.418, R = 0.342, P = 0.556, reflecting higher false negatives due to class imbalance. The same pattern can also be observed in Figure 5a and b, where the highlighted blue PR curve has the highest PR-AUC and seemingly fewer false negatives compared to the other models.
Additionally, as Table. I indicated, FM3M maintains strong results with few labels; at $N_{\text{train}}$ = 72, the larger pre-trained FM3M emerges as a better few-shot learner, achieving performance comparable to FM300K at $N_{\text{train}}$ = 729. These results underscore the advantages of large-scale SSL pre-training in managing imbalanced data and enhancing few-shot learning.
TABLE I: Performance on Imbalanced Loci Classification. The mean F1 scores, Recall (R), and Precision (P) from the five-fold cross-validation are presented. We compare performance of four model backbones: ViT-L trained from scratch, and DeepAndes pre-trained using 30K, 300K, and 3M images. $N_{\text{train}}$ represents the scale of “positive” images (containing loci) in the training dataset. The “positive” to “negative” ratio is 1:10 in both training and testing set. Entries marked with “–” indicate that the experiments do not converge or the values are not supported. For clarity, the highest metric values of each few-shot dataset evaluation are highlighted in bold.
@@ V-B Image Instance Retrieval
For loci image instance retrieval task, the performance of four frozen ViT-L backbones (Scratch, FM30K, FM300K, and FM3M) is compared in Table II. As demonstrated, for the image retrieval task, larger $k$ values result in lower mAP@k scores, indicating that retrieving more samples from the database also introduces more irrelevant ones relative to the query image class. It is also evident that pre-training improves retrieval performance. The Scratch model, without any pre-training, achieves the lowest mAP@k scores across all choices of k for this evaluation, with mAP@5 starting at 0.444. In contrast, pre-trained DeepAndes models—FM30K, FM300K, and FM3M—show progressively higher mAP values, with FM3M achieving the highest performance, with mAP@5 of 0.869. Figure 6 provides qualitative visualizations of the retrieved examples. As pre-training scales up, FM3M retrieves more relevant images (highlighted in blue boxes) that correctly match both the image class and features of the query image. In contrast, FM30K and the Scratch model retrieve incorrect image classes (highlighted in red boxes) among the top-5 retrieved examples. These results clearly demonstrate that SSL pre-training, particularly at a larger scale, enhances image-level feature representations and improves instance retrieval accuracy.
@@
TABLE II: Performance on Image Instance Retrieval. The mAP@$k$ (mean Average Precision within top-$k$ retrieved samples) for different choices of $k$ (5, 20, 50, and 100) is presented. Four frozen ViT-L backbones are compared: ViT-L (scratch) and DeepAndes pre-trained using 30K, 300K, and 3M images. For clarity, the highest metric values are highlighted in bold.
Figure 6: Examples of Retrieved Images from Different Pre-trained Models. The left column displays two example query images, with arrows pointing to key archaeological features. Both query images are from the positive class in the imbalanced loci classification dataset. The right columns show the top-5 retrieved images based on cosine similarity. The images with a red box indicate incorrect class retrieval, while the blue box highlights correct retrieval of both the image class and all archaeological features in the query image.
V-C Few-shot Loci Segmentation
Backbone Model Active Buildings1 Active Corrals2 Archaeological Corrals3 100% $N_{\text{train}}=10$ $N_{\text{train}}=20$ $N_{\text{train}}=30$ 100% $N_{\text{train}}=10$ $N_{\text{train}}=20$ $N_{\text{train}}=30$ 100% $N_{\text{train}}=10$ $N_{\text{train}}=20$ $N_{\text{train}}=30$ Frozen Scratch 29.1 8.2 21.3 20.6 38.3 11.6 14.6 21.0 8.4 FM30K 27.0 20.0 23.0 24.7 45.7 45.4 44.8 45.5 11.9 9.7 11.9 12.1 FM300K 55.2 41.4 47.3 51.4 60.9 55.5 56.5 59.3 27.8 19.0 17.8 24.8 FM3M 58.3 52.3 56.2 58.2 70.4 65.5 67.1 68.2 67.3 56.3 61.3 63.0 Finetuned Scratch 47.1 42.1 47.8 45.9 64.2 57.9 62.0 62.5 11.6 15.1 15.4 15.4 FM30K 53.8 44.0 52.8 53.2 66.4 61.3 63.7 64.8 17.7 16.2 15.6 18.4 FM300K 59.7 51.4 57.2 58.8 69.2 59.3 64.6 67.8 33.9 15.3 17.9 29.4 FM3M 69.3 57.8 63.7 66.5 81.1 70.8 73.5 75.5 84.8 51.3 72.5 81.0
@@ V-C Few-shot Loci Segmentation
TABLE III: Performance on Few-shot Segmentation of Three Loci Types. Mean DSC scores from five-fold cross-validation are presented, which include both transfer learning (frozen backbones) and fine-tuning (unfrozen backbones) across four models: Scratch, FM30K, FM300K, and FM3M. Specifically, $N_{\textbf{train}}$ indicates the scale of the training dataset. Bold entries denote the highest DSC score for each few-shot setting, while underscored entries indicate the second highest. Entries marked with “–” indicate that the experiments do not converge or the values are not supported.
For few-shot loci segmentation tasks, Table III summarizes the models’ performance on three loci types (i.e., active buildings, active corrals, and archaeological corrals). It is evident that as the pre-training scales increase—from 30K to 300K, and then to 3M—the model’s performance in both transfer learning (frozen) and fine-tuning improves compared to training from scratch. As highlighted in Table III, the best model performance across all few-shot datasets and tasks comes from the FM3M. For active/archaeological corrals segmentation tasks, simply training the linear segmentation head on top of the frozen FM3M backbone achieves the second highest DSC score across all experiments, demonstrating the benefits of our million-scale pre-training. On the other hand, models trained from scratch exhibit relatively low DSC scores, especially with small datasets. For example, at $N_{train}=10$, the DSC score for the Scratch model (frozen) on the Active Corrals dataset is only 11.6. In contrast, frozen pre-trained models show an improvement of over 30% starting from FM30K (45.4), with DSC scores continuing to rise as the pre-training scale increases. Although model performance improves with fine-tuning the entire ViT-L, transfer learning on the frozen FM3M backbone still exhibits DSC scores that are either comparable to or surpass those of models with smaller pre-training scales. This improvement is particularly noticeable when downstream data is very limited (e.g., $N_{train}=10$) across three segmentation tasks. Overall, these results highlight the effectiveness of million-scale DeepAndes in few-shot learning tasks, where the model can generalize well with limited labeled data. This makes it especially valuable in fields like archaeology, where data annotation is scarce.
@@
[…]
V-D DeepAndes Training and Fine-tuning Summary

### v2

@@ IV Data and Experiments
IV-A Data collection and Pre-processing
Data collection. The remote sensing satellite images used in this work were collected by the WorldView-2 and WorldView-3 satellites between 2010 and 2023, and imagery was provided by Maxar through the NextView Imagery License agreement with the National Geospatial-Intelligence Agency (NGA). We excluded images that are heavily obscured by clouds, as the underlying content is scarcely visible and therefore unsuitable for both machine learning analysis and manual survey. We observed that archaeological features—such as buildings, corrals, and related structures—are generally more discernible in imagery acquired during the dry season, whereas vegetation growth in the wet season often obscures these features. Consequently, whenever possible, we prioritize dry-season imagery (June through October). However, in cases where only wet-season data are available, we also include those images to ensure broader coverage.
Data pre-processing. First, the collected satellite images underwent color correction and orthographic correction using a coarse digital elevation model (DEM). The data are then pan-sharpened using the Bayesian fusion algorithm from Orfeo-Toolbox [21] to increase the spatial resolution of the multi-spectral imagery to 0.31 m for WorldView-2 imagery to match the native resolution of WorldView-3 imagery. In this work, all 8 spectral bands (four standard colors—red, green, blue, and near-infrared 1—and four new bands—coastal, yellow, red edge, and near-infrared 2) are used. Lastly, the imagery is re-sampled from 32 bits to 8 bits to reduce storage size and computational load.
@@ IV-B SSL with Different Pre-training Scales
As introduced in Section III-A, after data pre-processing, we construct the self-supervised pre-training dataset sourced from the entire Peruvian Andes. In total, the imagery covers survey regions of approximately 488,640 $km^{2}$ and includes 8 distinct land cover types. We densely sample 3 million image patches of size 256 256 from all surveyed regions to ensure a diverse, large-scale pre-training dataset.
In this work, we perform SSL using pre-training datasets of varying sizes to robustly evaluate the impact of DINOv2-based pre-training on downstream tasks. Specifically, we pre-train the DeepAndes backbone on datasets containing 30K, 300K, and 3 million image crops, denoted as FM30K, FM300K, and FM3M, respectively. Since our input data consists of 8-band imagery rather than standard RGB images, we also include a Scratch model—identical in architecture to the DeepAndes backbones (FM30K, FM300K, and FM3M) but initialized with random weights, i.e., without any pre-training—as a baseline for comparison.
All model backbones (FM30K, FM300K, FM3M, and Scratch) adopt the ViT-L/14 architecture, a Vision Transformer Large variant with 24 transformer layers, a hidden size of 1024, and a patch size of 14 14, totaling approximately 0.3 billion parameters. Input images are resized to 224 224 following the default configuration of the ViT-L/14 backbone in DINOv2. To accommodate 8-band imagery, the original patch embedding layer is replaced with a 2D convolutional layer that accepts 8 input channels. The ViT encoder outputs 257 tokens, including one class token and 256 patch tokens, each with a hidden dimension of 1,024.
@@ IV-C1 Imbalanced Loci Classification
Evaluation Metrics. Given the imbalanced nature of the classification task, we evaluate model performance using Precision, Recall, F1 score, and the area under the precision–recall curve (PR-AUC). Particularly, Precision is defined as the proportion of predicted positives that are correct. This is calculated as the ratio of true positives (TP) to the sum of TP and false positives (FP):
@@
The Recall (or sensitivity) measures the proportion of actual positives correctly identified. This is calculated as the ratio of TP to the the sum of TP and false negatives (FN):
@@
The F1 score combines both Precision and Recall into a single metric:
$F1\ score=2\times\frac{Precision\times Recall}{Precision+Recall}$
The PR-AUC summarizes performance across thresholds by integrating Precision as a function of Recall:
$PR\text{-}AUC=\int_{0}^{1}Precision(Recall)\,d(Recall)$
The combined use of these metrics enables a more comprehensive and fair evaluation in scenarios where the minority class (archaeological loci) is of interest.
@@ IV-C2 Image Instance Retrieval
Precision@i=Number of images retrieved at rank iiPrecision@i\ $\frac$$\text${Number of images retrieved at rank }i}{i}
@@
IV-C3 Few-shot Locus Segmentation
To evaluate the foundation model’s few-shot learning performance on pixel-level feature recognition, we focus on the semantic segmentation of three specific types of loci—active buildings, active corrals, and archaeological corrals. Three small-scale binary semantic segmentation datasets are constructed: the Active Buildings dataset contains 48 images, the Active Corrals dataset contains 55 images, and the Archaeological Corrals dataset contains 46 images. K-fold cross-validation (K=5) is employed for robust evaluation. The segmentation head is simple and directly utilizes the learned features from pre-training. Specifically, patch-level features are extracted from the images using patch tokens and concatenated with a simple one-layer linear segmentation head. Prior to the five-fold cross-validation experiments, we conducted several random seed trials to identify optimized hyperparameters for model training. Similar to the few-shot loci classification, we also evaluate the pre-trained model’s segmentation performance at different downstream dataset scales, including 10, 20, and 30 images.
@@ IV-C3 Few-shot Locus Segmentation
  • TPTP represents the number of pixels correctly classified as the loci in both the predicted and ground truth masks.
  • FPFP represents the number of pixels incorrectly classified as loci in the predicted mask but not in the ground truth mask.
  • FNFN represents the number of pixels incorrectly classified as background in the predicted mask but actually belonging to the loci in the ground truth mask.
@@ V Results
TABLE II: Performance on Imbalanced Loci Classification. The mean F1 scores, Recall (R), and Precision (P) from the five-fold cross-validation are presented. We compare performance of four model backbones: ViT-L trained from scratch, and DeepAndes pre-trained using 30K, 300K, and 3M images. $N_{\text{train}}$ represents the scale of “positive” images (containing loci) in the training dataset. The “positive” to “negative” ratio is 1:10 in both training and testing set. Entries marked with “–” indicate that the experiments do not converge or the values are not supported. For clarity, the highest metric values of each few-shot dataset evaluation are highlighted in bold.
@@ V-A Imbalanced Loci Classification
Both Table II and Figure 5 display the five-fold cross-validation results for imbalanced loci classification. Specifically, Table II details the mean F1 scores, Precision (Prec), and Recall (Rec). Figure 5a illustrates the Precision-Recall (PR) curves, while Figure 5b presents the confusion matrices for each model, highlighted using a representative hold-out fold.
As shown in Table II, the Scratch model struggles with limited and imbalanced data, converging only at $N_{train}=729$ (F1 = 0.544) and $N_{train}=365$ (F1 = 0.402), while failing entirely at smaller datasets. In contrast, DeepAndes models with SSL pre-training improve consistently across all dataset sizes. At $N_{train}=72$, FM3M achieves F1 = 0.83, Recall = 0.825, and Precision = 0.837, effectively balancing false positives and false negatives. Smaller pre-trained models decline in performance, with FM300K dropping to F1 = 0.728, Rec = 0.671, Prec = 0.812 and FM30K to F1 = 0.418, Rec = 0.342, Prec = 0.556, reflecting higher false negatives due to class imbalance. The same pattern can also be observed in Figure 5a and b, where the highlighted blue PR curve has the highest PR-AUC and seemingly fewer false negatives compared to the other models.
Additionally, as Table. II indicated, FM3M maintains strong results with few labels; at $N_{train}=72$, the larger pre-trained FM3M emerges as a better few-shot learner, achieving performance comparable to FM300K at $N_{train}=729$. These results underscore the advantages of large-scale SSL pre-training in managing imbalanced data and enhancing few-shot learning.
@@ V-B Image Instance Retrieval
For loci image instance retrieval task, the performance of four frozen ViT-L backbones (Scratch, FM30K, FM300K, and FM3M) is compared in Table III. As demonstrated, for the image retrieval task, larger $k$ values result in lower mAP@k scores, indicating that retrieving more samples from the database also introduces more irrelevant ones relative to the query image class. It is also evident that pre-training improves retrieval performance. The Scratch model, without any pre-training, achieves the lowest mAP@k scores across all choices of k for this evaluation, with mAP@5 starting at 0.444. In contrast, pre-trained DeepAndes models—FM30K, FM300K, and FM3M—show progressively higher mAP values, with FM3M achieving the highest performance, with mAP@5 of 0.869. Figure 6 provides qualitative visualizations of the retrieved examples. As pre-training scales up, FM3M retrieves more relevant images (highlighted in blue boxes) that correctly match both the image class and features of the query image. In contrast, FM30K and the Scratch model retrieve incorrect image classes (highlighted in red boxes) among the top-5 retrieved examples. These results clearly demonstrate that SSL pre-training, particularly at a larger scale, enhances image-level feature representations and improves instance retrieval accuracy.
@@
TABLE III: Performance on Image Instance Retrieval. The mAP@$k$ (mean Average Precision within top-$k$ retrieved samples) for different choices of $k$ (5, 20, 50, and 100) is presented. Four frozen ViT-L backbones are compared: ViT-L (scratch) and DeepAndes pre-trained using 30K, 300K, and 3M images. For clarity, the highest metric values are highlighted in bold.
Figure 6: Examples of Retrieved Images from Different Pre-trained Models. The left column displays two example query images, with arrows pointing to key archaeological features. Both query images are from the positive class in the imbalanced loci classification dataset. The right column presents the top-5 retrieved images based on cosine similarity. The images with a red box indicate incorrect class retrieval, while the blue box highlights correct retrieval of both the image class and all archaeological features in the query image.
V-C Few-shot Locus Segmentation
Backbone Model Active Buildings1 Active Corrals2 Archaeol. Corrals3 100% $N_{\text{train}}=10$ $N_{\text{train}}=20$ $N_{\text{train}}=30$ 100% $N_{\text{train}}=10$ $N_{\text{train}}=20$ $N_{\text{train}}=30$ 100% $N_{\text{train}}=10$ $N_{\text{train}}=20$ $N_{\text{train}}=30$ Frozen Scratch 29.1 8.2 21.3 20.6 38.3 11.6 14.6 21.0 8.4 FM30K 27.0 20.0 23.0 24.7 45.7 45.4 44.8 45.5 11.9 9.7 11.9 12.1 FM300K 55.2 41.4 47.3 51.4 60.9 55.5 56.5 59.3 27.8 19.0 17.8 24.8 FM3M 58.3 52.3 56.2 58.2 70.4 65.5 67.1 68.2 67.3 56.3 61.3 63.0 Finetuned Scratch 47.1 42.1 47.8 45.9 64.2 57.9 62.0 62.5 11.6 15.1 15.4 15.4 FM30K 53.8 44.0 52.8 53.2 66.4 61.3 63.7 64.8 17.7 16.2 15.6 18.4 FM300K 59.7 51.4 57.2 58.8 69.2 59.3 64.6 67.8 33.9 15.3 17.9 29.4 FM3M 69.3 57.8 63.7 66.5 81.1 70.8 73.5 75.5 84.8 51.3 72.5 81.0
@@ V-C Few-shot Locus Segmentation
TABLE IV: Performance on Few-shot Segmentation of Three Loci Types. Mean DSC scores from five-fold cross-validation are presented, which include both transfer learning (frozen backbones) and fine-tuning (unfrozen backbones) across four models: Scratch, FM30K, FM300K, and FM3M. Specifically, $N_{\textbf{train}}$ indicates the scale of the training dataset. Bold entries denote the highest DSC score for each few-shot setting, while underscored entries indicate the second highest. Entries marked with “–” indicate that the experiments do not converge or the values are not supported.
For few-shot locus segmentation, Table IV summarizes the models’ performance on three locus types (i.e., active buildings, active corrals, and archaeological corrals). It is evident that as the pre-training scales increase—from 30K to 300K, and then to 3M—the model’s performance in both transfer learning (frozen) and fine-tuning improves compared to training from scratch. As highlighted in Table IV, the best model performance across all few-shot datasets and tasks comes from the FM3M. For active/archaeological corrals segmentation tasks, simply training the linear segmentation head on top of the frozen FM3M backbone achieves the second highest DSC score across all experiments, demonstrating the benefits of our million-scale pre-training. On the other hand, models trained from scratch exhibit relatively low DSC scores, especially with small datasets. For example, at $N_{train}=10$, the DSC score for the Scratch model (frozen) on the Active Corrals dataset is only 11.6. In contrast, frozen pre-trained models show an improvement of over 30% starting from FM30K (45.4), with DSC scores continuing to rise as the pre-training scale increases. Although model performance improves with fine-tuning the entire ViT-L, transfer learning on the frozen FM3M backbone still exhibits DSC scores that are either comparable to or surpass those of models with smaller pre-training scales. This improvement is particularly noticeable when downstream data is very limited (e.g., $N_{train}=10$) across three segmentation tasks. Overall, these results highlight the effectiveness of million-scale DeepAndes in few-shot learning tasks, where the model can generalize well with limited labeled data. This makes it especially valuable in fields like archaeology, where data annotation is scarce.
@@
Additionally, we present the qualitative visualizations, paired with the Table IV results, for the three types of loci segmentation tasks in Figure 7. An example patch and its segmentation results from different models are shown. As shown, segmenting archaeological corrals (bottom panel) is the most challenging, as there is little to no difference between the inside and outside, except for the boundary pixels. Segmenting active corrals (top panel) is the easiest, as visual residues of active use (e.g., removal of vegetation through grazing) creates high contrast between the interior and exterior of corrals. Active building segmentation (middle panel) shows medium performance generally. The variability in building colors and their small size makes it difficult for the model to converge. As shown, under few-shot settings ($N_{train}=10,\ 20,\ 30$), Scratch and FM30K struggled to achieve reliable performance, as evidenced by the transfer learning results for even the best-performing task (top panel).
To some extent, fine-tuning the entire model, though more costly, helps address these issues. For example, FT-FM3M, which uses the DeepAndes (3M) backbone, effectively captures foreground pixels and, with a simple linear segmentation head, achieves the highest segmentation quality across the three tasks. However, in active building segmentation (middle panel), transfer learning with FM3M still achieves strong performance, outperforming other frozen backbones and demonstrating results comparable to the more costly full-model fine-tuning approach (e.g., FT-FM300K). For the most challenging task (bottom panel), archaeological corral segmentation, FM3M and FT-FM3M clearly outperform other models, which struggle to capture the foreground pixels and produce many false positives.
V-D Ablation Study
V-D1 Performance Comparison
As shown in Table V, we benchmark our DeepAndes backbone (FM3M) against several representative baselines: a randomly initialized scratch model, SSL backbones including MAE and MoCo-V2, and SatMAE, a domain-specific remote sensing foundation model. Performance is evaluated across three selected tasks introduced previously—zero-shot image retrieval (Top-5 and Top-100 mAP), few-shot classification (F1, Recall, and Precision), and few-shot segmentation (DSC). For classification and segmentation, we report results for both the full training set and a severely constrained few-shot setting ($N_{train}=72$ for classification and $N_{train}=10$ for segmentation) to simulate real-world low-resource conditions. The best performance for each metric is highlighted.
Across all tasks, DeepAndes consistently outperforms the baselines, highlighting the benefits of large-scale DINOv2-based pre-training on 8-band imagery. In zero-shot image retrieval, it achieves the highest Top-5 and Top-100 mAP scores (0.869 and 0.69), showing the strong generalization capability of DeepAndes without task-specific fine-tuning.
In the few-shot classification task, models with SSL pre-training outperform the scratch model, maintaining relatively high F1, precision, and recall scores. Notably, the other remote sensing foundation model, such as SatMAE, performs worse than DeepAndes, suggesting that pre-training on different satellite sources introduces a domain gap. DeepAndes maintains a strong precision-recall balance across settings. While MAE and MoCo-V2 perform comparably in the high-data regime, their performance declines more noticeably under the limited supervision ($N_{train}=72$).
For few-shot segmentation, DeepAndes delivers the top DSC scores in 4 out of 6 settings, specifically for first two archaeological features (active buildings and corrals). Its performance remains more consistent at $N_{train}=10$, highlighting its robustness in low-data regimes. However, it is worth noting that for the segmentation of archaeological corrals, DeepAndes ranks second to MAE (0.71 vs. 0.67 and 0.66 vs. 0.56), suggesting that incorporating MAE-like strategies during pre-training may enhance the model’s ability to capture dense features.
Method Zero-shot Retrieval Few-shot Classification Few-shot Segmentation (DSC) Top-5 Top-100 $N_{\text{train}}=729$ $N_{\text{train}}=72$ Active Buildings Active Corrals Archaeol. Corrals mAP ↑ mAP ↑ F1 ↑ Rec ↑ Prec ↑ F1 ↑ Rec ↑ Prec ↑ 100% $N_{\text{train}}=10$ 100% $N_{\text{train}}=10$ 100% $N_{\text{train}}=10$ Scratch 0.444 0.296 0.544 0.457 0.735 0.291 0.082 0.383 0.116 0.084 MoCo-V2 0.511 0.343 0.88 0.902 0.864 0.78 0.727 0.831 0.184 0.032 0.559 0.463 0.328 0.307 MAE 0.709 0.473 0.92 0.901 0.937 0.82 0.831 0.804 0.458 0.20 0.673 0.298 0.71 0.66 SatMAE 0.578 0.371 0.87 0.885 0.853 0.71 0.70 0.711 0.051 0.04 0.163 0.177 0.026 0.04 DeepAndes 0.869 0.69 0.886 0.876 0.894 0.83 0.825 0.837 0.583 0.523 0.704 0.654 0.673 0.563
TABLE V: We benchmark DeepAndes (FM3M) against additional representative baselines: a randomly initialized Scratch model, self-supervised backbones including MoCo-V2, MAE, and SatMAE, a domain-specific remote-sensing foundation model. Models are evaluated across three selected tasks: zero-shot image retrieval (Top-5 and Top-50 mAP), few-shot classification (F1, Recall, and Precision), and few-shot segmentation (DSC). For segmentation, the backbone is frozen and only a linear segmentation head is trained. Few-shot results are reported for both the full training set and the most constrained subset ($N_{\text{train}}=72$ for classification and $N_{\text{train}}=10$ for segmentation) to simulate data-limited conditions. The highest value for each evaluation metric is highlighted in bold.
V-D2 Computational Complexity
Table VI reports the computational complexity of representative backbone models. Latency reflects the inference speed, measured as milliseconds per sample (ms/sample) on an NVIDIA RTX A6000 GPU with batch size 1 and input resolution of 224 224. Each value is averaged over 100 runs. FLOPs are reported in units of $1\times 10^{9}$ operations (GFLOPs), and the number of parameters is given in millions (M). Transformer-based models (MAE, SatMAE, and DeepAndes) incur higher inference costs than the ResNet50 baseline (MoCo-V2). DeepAndes shows the highest FLOPs and latency, reflecting a trade-off between computational efficiency and downstream performance. This suggests further improvements on model efficiency—through compression or lightweight designs—while preserving generalization.
Method Backbone Latency (ms) FLOPs (G) Params (M)
MoCo-V2 ResNet-50 2.61 8.65 23
MAE ViT-L/16 7.55 84.15 304
SatMAE ViT-L/16 7.83 84.15 330
DeepAndes ViT-L/14 10.69 156.29 304
TABLE VI: Computational Complexity Comparison. Latency (ms/sample) is measured on an NVIDIA RTX A6000 GPU with batch size 1 and input resolution of $224\times 224$. Each value is averaged over 100 runs. FLOPs are reported in $1\times 10^{9}$ operations (GFLOPs), and parameters in millions (M).
V-E DeepAndes Training and Fine-tuning Summary